# SHaRe-SSM: An Oscillatory Spiking Neural Network for Target Variable Modeling in Long Sequences

### v1

@@ A.1.1 Implicit Discretization (IM):
This formulation highlights the stabilizing effect of the implicit method: the matrix $S=(I+\Delta t^{2}\Omega)^{-1}$ is a Schur complement that acts as a preconditioner that suppresses high-frequency components. Consequently, the eigenvalues of $M^{\text{IM}}$ remain bounded within the unit circle for a wide range of $\Delta t$, leading to improved numerical stability. The Schur complement can be computed in $\mathcal{O}(m)$ instead of the typical $\mathcal{O}(m^{3})$ operations using Gauss-Jordan elimination.
@@
A.1.2 Implicit-Explicit Discretization (IMEX):
We also utilize an implicit-explicit (IMEX) scheme for discretizing the second-order harmonic oscillator system, similar to [21]. IMEX methods treat the stiff terms implicitly and the non-stiff or input terms explicitly, resulting in a balanced scheme that enables stable yet undamped oscillations. As shown in [21], such schemes preserve the total energy of the system and therefore are particularly well-suited for learning long-range sequential patterns without introducing artificial dissipation.
The update equations under the IMEX discretization are given by:
@@ A.1.2 Implicit-Explicit Discretization (IMEX):
We analyze the characteristic polynomial:

### v2

@@ A.1.1 Implicit discretisation (IM):
This formulation highlights the stabilising effect of the implicit method: the matrix $S=(I+\Delta t^{2}\Omega)^{-1}$ is a Schur complement that acts as a preconditioner, suppressing high-frequency components. Consequently, the eigenvalues of $M^{\text{IM}}$ remain bounded within the unit circle for a wide range of $\Delta t$, leading to improved numerical stability. The Schur complement can be computed in $\mathcal{O}(m)$ instead of the typical $\mathcal{O}(m^{3})$ operations using Gauss-Jordan elimination.
Proposition A.1 is rephrased from proposition 3.1 of Rusch and Rus. [rusch2025oscillatory]
@@
Proposition A.2 is adapted from proposition 3.2 of Rusch and Rus. [rusch2025oscillatory]
@@
A.1.2 Implicit-Explicit discretisation (IMEX):
We also utilise an implicit-explicit (IMEX) scheme for discretising the second-order harmonic oscillator system, similar to rusch2025oscillatory. IMEX methods treat the stiff terms implicitly and the non-stiff or input terms explicitly, resulting in a balanced scheme that enables stable yet undamped oscillations. As shown in rusch2025oscillatory, such schemes preserve the total energy of the system and therefore are particularly well-suited for learning long-range sequential patterns without introducing artificial dissipation.
The update equations under the IMEX discretisation are given by:
@@ A.1.2 Implicit-Explicit discretisation (IMEX):
Proposition A.3 is adapted from proposition E.1 of Rusch and Rus. [rusch2025oscillatory]
@@
We analyse the characteristic polynomial:
@@
A.2 Parallel scan in SHaRe-SSM
Parallel scans [kogge1973parallel, blelloch1990prefix] exploit associativity to reduce recurrent computation from $\mathcal{O}(N)$ to $\mathcal{O}(\log N)$. Originally developed for RNNs, they have recently been adapted to state-space models [s5], enabling efficient architectures such as LRUs [lru] and Mamba [mamba]. In our setting, parallel scans accelerate linear updates, with spike functions applied afterwards. Following rusch2025oscillatory, we define an associative binary operation:
$(a_{1},a_{2})\bullet(b_{1},b_{2})=(b_{1}\cdot a_{1},\;b_{1}\cdot a_{2}+b_{2}),$ (7)
where $\cdot$ denotes matrix-matrix or matrix-vector multiplication. Applying a parallel scan to the input sequence $\{(M,F_{n})\}$ efficiently solves
$s_{n}=Ms_{n-1}+F_{n},$ (8)
with the second tuple element storing $x_{n}$. Efficiency is achieved by exploiting structured matrices (e.g., diagonal block $2\times 2$ forms in $M_{IM}$ and $M_{IMEX}$), where each multiplication is linear in the hidden dimension. We use this formulation to implement IM and IMEX discretisations for HRF neurons. Algorithm 1 summarises the SHaRe-SSM implementation.
A.3 Energy Computation
We compute the energy for a LinOSS block and compare it to our SHaRe-SSM block. We can observe that our block doesn’t perform any matrix multiplications and is well-suited for neuromorphic hardware. For event-based sensors, we can detach the encoder head and feed data directly to the model for real-time sequential processing. Spike rates from linear layers, SHaRe-SSM neuron and post weighted spike-mixing layers are given by $f^{\theta},f^{\theta_{C}},f^{\theta_{D}}$ for Sequence Length(L), State Size(P), Hidden Size(H), respectively.
The ratio of Energy consumed by LinOSS/SHaRe-SSM computed by:
$\text{Ratio}=\frac{E_{MAC}\times N\left({2LPH}+{(7+2)LH^{2}}\right)}{E_{AC}\sum_{i=1}^{N}\left({\left(\sum_{j=1}^{i}f^{\theta}_{j}+f^{\theta_{C}}_{i}\right)LPH}_{B,C}+{f_{i}^{\theta_{D}}LH^{2}}\right)}$
We estimate the theoretical energy consumption of our model based on prior works [spikingssm, pspikessm, wu2025spikf]. Accordingly, we assume that MAC and AC operations are implemented on a 45nm hardware [horowitz20141], where $E_{\text{MAC}}=4.6\text{pJ}$ and $E_{\text{AC}}=0.9\text{pJ}$. Notably, as discussed in  spikingssm, the computational cost of multiplication of a floating-point weight by a binary activation number is represented as an addition-only operation. For ANNs, the theoretical energy consumption of a block $n$ is given by $4.6\text{pJ}\times\text{FLOPs}(n)$. For SNNs, the energy consumption for $n$ is given by $0.9\text{pJ}\times\text{SOPs}(n)$. Calculating theoretical energy consumption requires first calculating the synaptic operations, $\text{SOPs}(n)=f_{r}\times\text{FLOPs}(n)$, $f_{r}$ is the firing rate of the input spike train of the block/layer, $\text{FLOPs}(n)$ refers to the number of floating-point operations in layer $n$, equivalent to the number of multiply-and-accumulate (MAC) operations. SOPs denote the number of spike-based accumulate (AC) operations.
Since $\Omega$ is a diagonal matrix, it can be efficiently implemented using parallel scans with computations of order $\mathcal{O}(P\,\log(L))$ (Total computations are $\mathcal{O}(PL)$), which is negligible. Also, D is $\mathcal{O}(HL)$, which is much smaller than $\mathcal{O}(LPH)$, i.e., computation for B & C matrices or even $\mathcal{O}(LH^{2})$, which is for the GLU layer. [rusch2025oscillatory] uses GeLU and GLU non-linearities, which we replace with a linear layer. And, as described in [yu2023metaformer], GeLU consumes 14 FLOPs per operation. Moreover, GLU has twice as many FLOPs as a linear layer.
Table 3: Hyperparameters for the Best model for each dataset
Dataset Method LR Hidden State Blocks Time Kernel
UCI-HAR IM 1e-3 128 256 2 False -
UCI-HAR IMEX 1e-3 128 256 2 False -
SHAR IM 1e-3 128 256 2 False -
SHAR IMEX 1e-3 128 256 2 False -
Worms IM 1e-4 128 256 2 False -
Worms IMEX 1e-3 128 64 2 False -
PPG IM 1e-3 64 64 6 False 8
PPG IMEX 1e-3 128 256 6 False 16